[Brahms-l] Shift report 20050125 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Jan 25 2005 - 23:57:04 EST
There was a Quench Link Interlock at shift start.

Beam achieved at about 17:30
Runs with FS 8 deg 1723B, MRS 90 deg B2255:
          tr2       tr3
12490   50k          44k
12491   67k          60k
12492   60k          55k
12493   39k          35k
12494   30k          27k
12495  TPC Pedestal run
12496  CAMAC Pedestal run

This brings the total FS 8 deg 1723B has achieved the total 600k 
when adding the previous runs.  Trig 3 is now about 1.2M (200k in
this shift)

Detector problems.
Noticed MCh19 of TOF-II was off and could not be turned on.
Tile # 22 and 23 have no ADC signals (see e-log)
T3-3-v2 has wires missing

Perhaps all of these are known problems, but thought to mention them
anyway.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Jan 25 23:57:25 2005

This archive was generated by hypermail 2.1.8 : Tue Jan 25 2005 - 23:57:48 EST